OpenStudy (decarr432):

Given that -1,-2 is on the graph of f(x), find the corresponding point for the function f(x)+4

OpenStudy (anonymous):

looks like you add 4 to the output right?

OpenStudy (decarr432):

I think so but i'm not entirely sure

OpenStudy (anonymous):

for example if \(f(2)=8\) then \(f(2)+4=8+4=12\)

OpenStudy (anonymous):

so add 4 to the second coordinate, that is all

OpenStudy (decarr432):

so -1,2

OpenStudy (anonymous):

yes